# Impact of flavour coupling on $SO(10)$-inspired leptogenesis

### v1

@@ 3 Including flavour coupling
The expression in Eq. (27) for the final $B-L$ asymmetry assumes that the evolution of the three flavoured asymmetries $\Delta_{\alpha}\equiv B/3-L_{\alpha}$ ($\alpha=e,\mu,\tau$), where the $L_{\alpha}$’s are the lepton asymmetries in lepton doublets, proceeds independently of each other, in an uncoupled way. Sphaleron process conserve the $\Delta_{\alpha}$’s and the total $B-L$ asymmetry is the sum of the three. In this way one can solve the set of uncoupled differential equations for the $\Delta_{\alpha}$’s and at the end $N_{B-L}^{\rm f}=\sum_{\alpha}N_{\Delta_{\alpha}}^{\rm f}$ and the result is the expression in Eq. (27) [19, 51, 36, 52]. The assumption of uncoupled evolution of the flavoured asymmetries relies on a simple picture where: $N_{2}$-decays source initially the asymmetry in the flavour direction ${\ell}_{2}$; decoherence of leptons states results into an independent washout from $N_{2}$-inverse processes for the two flavoured asymmetries $\Delta_{\tau}$ and $\Delta_{e}+\Delta_{\mu}$ at $T\sim M_{2}$ (two flavour regime) and into an independent washout from $N_{1}$-inverse processes for all three $\Delta_{\alpha}$ at $T\sim M_{1}$; finally, it assumes that the washout is just depending on each $\Delta_{\alpha}$ separately. Within this simple picture, though decoherence splits the evolution of the asymmetry first in two and then in three flavoured asymmetries, these still develop independently of each other because the the washout in a certain flavour is always depending, linearly in very good approximation, just on the asymmetry in that flavour. However, this picture is incomplete because it neglects spectator processes of different nature [34, 35, 17, 51]: (i) the generation of an asymmetry from $N_{2}$-decays into lepton doublets is also accompanied by the generation of a hypercharge asymmetry into the Higgs bosons and (ii) by a baryonic asymmetry into squarks via sphaleron processes; (iii) the lepton asymmetry from lepton doublets is also redistributed to RH charged particles. In this way the washout of a $\Delta_{\alpha}$ asymmetry also depends on the asymmetries stored into the Higgs doublets (this is the primary source of flavour coupling) and into quarks (this a secondary source of flavour coupling). To be more clear, since the Higgs doublet asymmetry is flavour blind, if a $\Delta_{\alpha}$ asymmetry has been generated in a certain flavour, necessarily at the washout the asymmetry into the Higgs doublets will induce the generation of a $\Delta_{\beta\neq\alpha}$ asymmetry in the other flavour(s) $\beta$: therefore, inverse processes do in this case generate a flavoured asymmetry rather than wash it out. This means that the washout strictly speaking refers to the Higgs asymmetry and the $N_{2}$-decay produced $\Delta_{\alpha}$ asymmetry but not to $\Delta_{\beta\neq\alpha}$. This is the primary source of flavour coupling. Secondarily, the redistribution of the asymmetries into quarks and RH charged particles also couples the $\Delta_{\alpha}$ asymmetries.
@@ 3.1 A more tangled expression for the final asymmetry
Solving the system of coupled differential equations at the production, for $T\sim M_{2}$, and at the $N_{1}$-washout, for $T\sim M_{1}$, leads to the following analytical expression for each final $\Delta_{\alpha}$ asymmetry [36]:
@@
Before explaining in detail the meaning and definition of all involved quantities in this general expression, notice that now the final $B-L$ asymmetry, though still given by $N^{\rm f}_{B-L}=\sum_{\alpha=e,\mu,\tau}N^{\rm f}_{\Delta_{\alpha}}$, is the sum of 27 terms instead of just 3. The matrix
@@
is the matrix that diagonalizes $P^{0}_{1}$, i.e. $V\,P^{0}_{1}\,V^{-1}=P^{0}_{1^{\prime\prime}}\equiv{\rm diag}(P^{0}_{1e^{\prime\prime}},P^{0}_{1\mu^{\prime\prime}},P^{0}_{1\tau^{\prime\prime}})$, where the matrix $P^{0}_{1}$ is defined as
@@
and the three-flavour coupling matrix is given by [54]
$C^{(3)}\equiv\left(\begin{array}[]{ccc}C_{ee}^{(3)}&C_{e\mu}^{(3)}&C_{e\tau}^{(3)}\\ C_{\mu e}^{(3)}&C_{\mu\mu}^{(3)}&C_{\mu\tau}^{(3)}\\ C_{\tau e}^{(3)}&C_{\tau\mu}^{(3)}&C_{\tau\tau}^{(3)}\end{array}\right)=\left(\begin{array}[]{ccc}188/179&32/179&32/179\\ 49/358&500/537&142/537\\ 49/358&142/537&500/537\end{array}\right)\,,$ (43)
and we defined $P^{0}_{1\alpha}\equiv K_{1\alpha}/K_{1}$. The rotated flavoured decay parameters are defined as $K_{1\alpha^{\prime\prime}}\equiv P^{0}_{1\alpha^{\prime\prime}}\,K_{1}$.
@@
i.e. $U\,P^{0}_{2}\,U^{-1}={\rm diag}(P^{0}_{2\tau^{\bot^{\prime}}_{2}},P^{0}_{2\tau^{\prime}})$, where the two-flavour coupling matrix is given by [54]
@@
and we defined $P^{0}_{2\alpha}\equiv K_{2\alpha}/K_{2}$. We have now fully defined all quantities entering the expression for the flavoured asymmetries in Eq. (39). As we said, the final $B-L$ asymmetry is still obtained as the sum of the three flavoured asymmetries and it can be checked that, for $U=V=I$, one recovers the expression Eq. (27) valid when flavour coupling is neglected. We can also unpack the first sum in Eq. (39) and write explicitly:
@@
This expression shows how now each flavoured asymmetry $N^{\rm f}_{\Delta_{\alpha}}$ is not simply given by one term containing a $N_{1}$ washout exponential suppression term described by $e^{-3\pi\,K_{1\alpha}/8}$ but it also contains terms that are washed out by exponentials $e^{-3\pi\,K_{1\delta^{\prime\prime}}/8}$ with $\delta\neq\alpha$. In this way, even though $K_{1\alpha}\gg 1$, there can still be unsuppressed contributions to $N^{\rm f}_{\Delta_{\alpha}}$ from terms with $K_{1\delta}\ll 1$. Even though these terms are weighted by factors $V^{-1}_{\alpha\delta}$, containing off-diagonal terms of the $C^{(3)}$ matrix that are ${\cal O}(0.1)$, they can be dominant in some cases and therefore, in general, they have to be accounted for.
@@ 3.2 Successful leptogenesis solutions
Despite the much higher intricacy of the expression Eq. (39), accounting for flavour coupling, compared to Eq. (27), neglecting flavour coupling, the discussion we made in 2.3 on the parameter dependence and the consequent successful leptogenesis condition Eq. (38) still holds. Flavour coupling does not introduce any new parameter. It only introduces many new terms in the expression for the final asymmetry that are proportional to the off-diagonal terms in the flavour coupling matrix. These are usually sub-dominant and give a small correction but for particular values of the parameters they can yield new solutions in regions of parameter space that would otherwise be unaccessible when flavour coupling is neglected.
The search of these solutions is the main objective of including flavour coupling. The result is shown in the right panel of Fig. 1. Here we show again the three-dimensional projection of a scatter plot of a about $2\times 10^{6}$ solutions obtained for $\alpha_{2}=5$ and scanning over the 15 parameters in $U$ and $V_{L}$, respecting the experimental constraints in Eq. (6). The solutions have been again obtained imposing $\chi^{2}<\chi^{2}_{\rm max}=25$ and $\eta_{B}>6.01\times 10^{-10}$. The first thing one can notice is that the main structure of the allowed region is not changed by flavour coupling and, in particular, one still has a lower bound on $m_{1}$. On the other hand, it can be noticed how there are new muonic solution appearing in a region that was before excluded, at values of $m_{1}\sim 10\,{\rm meV}$. Moreover, electronic solutions, indicated by red points, now appear in the region of $\tau_{A}$ solutions.
@@ 3.2.1 Two new types of solutions
Let us have a closer look a these new solutions. As we did in the last section for the three types of old solutions, in Fig. 9 we plot different quantities ($M_{I},\varepsilon_{\alpha},K_{1\alpha},K_{2\tau},K_{2\tau^{\bot}_{2}}$) versus $m_{1}$ for two benchmark new solutions, one muonic and one electronic, with values of the parameters given in Table 2.
@@
As one can see, successful leptogenesis is satisfied for $m_{1}=42\,{\rm meV}$ and $m_{1}=2.1\,{\rm meV}$ for the muon-dominated and the electron-dominated solution, respectively. One can also see how the dominant $C\!P$ asymmetry is the tauonic $C\!P$ asymmetry while both the muonic and the electronic ones would be too small, in the absence of flavour coupling, to yield the observed baryon asymmetry. Therefore, the existence of these two new types of solutions crucially relies on the account of flavour coupling (without flavour coupling the muon and electron asymmetries, respectively, would be orders of magnitude below the observed value). We should add, however, that these new electronic solutions are very marginal ones. This can be somehow understood by the fact that the peak of the asymmetry in the shown benchmark case is just above the observed value. In the $2\times 10^{6}$ point scatter plot we generated, we obtained only less than $200$ electronic solution (i.e., less than $0.1\%$). The new muonic solutions are less special and we obtained about 2k of them (i.e., $\sim 1\%$).
@@ 3.2.2 Lower bound on $m_{1}$
In the right panel of Fig. 3 we again show the isocontour lines for the $m_{1}$ lower bound in the $\delta$ versus $\theta_{23}$ plane. One can see that they are just slightly modified by the inclusion of flavour coupling and, for example, the absolute minimum value gets slightly relaxed from $m_{1}\geq 0.89\,{\rm meV}$ to $m_{1}\geq 65\,{\rm meV}$. The most remarkable difference is that now the excluded region by the cosmological upper bound Eq. (2), in red in Fig. 3, gets reduced and first octant values of $\theta_{23}$ are now possible also for values $\delta\sim\pi$. For these values of $\delta$, one has that values of the atmospheric mixing angle $\theta_{23}\gtrsim 47^{\circ}$ are still excluded even when flavour coupling is included. In particular, one can see how the best fit point falls now in the green region that is the one opened by the muonic solutions induced by the account of flavour coupling.
@@ 3.2.3 Upper bound on $\theta_{23}$
[…]
@@ 3.2.4 Excluded regions for $\delta$
In the right panel of Fig. 5 we show the constraints in the plane $\delta$ versus $m_{1}$. One can see how the new muonic solutions tend to fill some existing gaps for $m_{1}\simeq 10\,{\rm meV}$ and $\delta$ in the range between $\pi/2$ and $23\pi/$ 2. However, the hole at $\delta\simeq\pi/2$, though shrinking, still survives. For $m_{1}$ much below 10 meV the constraints are practically unchanged.

### v2

@@ 3 Including flavour coupling
The expression for the final BLB-L asymmetry in Eq. (27) assumes that the evolution of the three flavoured asymmetries $\Delta_{\alpha}\equiv B/3-L_{\alpha}$ ($\alpha=e,\mu,\tau$), where the $L_{\alpha}$’s are the lepton asymmetries in lepton doublets, proceeds independently of each other, in an uncoupled way. Sphaleron process conserve each $\Delta_{\alpha}$ and the total BLB-L asymmetry is the sum of the three. In this way one can solve the set of uncoupled differential equations for the $\Delta_{\alpha}$’s. At the end $N_{B-L}^{\rm f}=\sum_{\alpha}N_{\Delta_{\alpha}}^{\rm f}$ and the result is the expression in Eq. (27) vives , bounds , fuller , density . The assumption of uncoupled evolution of the flavoured asymmetries relies on a simple picture where: $N_{2}$-decays source initially the BLB-L asymmetry in a specific direction in flavour space (a ray, in space vector language) denoted by ${\ell}_{2}$; decoherence of leptons states results into an independent washout from $N_{2}$-inverse processes for the two flavoured asymmetries $\Delta_{\tau}$ and $\Delta_{e}+\Delta_{\mu}$ at $T\sim M_{2}$ (two flavour regime) and into an independent washout from $N_{1}$-inverse processes for all three $\Delta_{\alpha}$ at $T\sim M_{1}$; finally, it assumes that the washout is just depending on each $\Delta_{\alpha}$ separately. Within this simple picture, though decoherence splits the evolution of the asymmetry first in two and then in three flavoured asymmetries, these still develop independently of each other because the the washout in a certain flavour is always depending, linearly in very good approximation, just on the asymmetry in that flavour. However, this picture is incomplete because it neglects spectator processes of different nature Barbieri:1999ma , Buchmuller:2001sr , flavoureffects , bounds : (i) the generation of an asymmetry from $N_{2}$-decays into lepton doublets is also accompanied by the generation of a hypercharge asymmetry into the Higgs bosons and (ii) by a baryonic asymmetry into quarks via sphaleron processes; (iii) the lepton asymmetry from lepton doublets is also redistributed to RH charged particles. In this way the washout of a $\Delta_{\alpha}$ asymmetry also depends on the asymmetries stored into the Higgs doublets (this is the primary source of flavour coupling) and quarks (this a secondary source of flavour coupling). To be more specific, since the Higgs doublet asymmetry is flavour blind, if a $\Delta_{\alpha}$ asymmetry has been generated in a certain flavour, necessarily at the washout the asymmetry into the Higgs doublets will induce the generation of a $\Delta_{\beta\neq\alpha}$ asymmetry in the other flavour(s) $\beta$: therefore, inverse processes do in this case generate a flavoured asymmetry rather than wash it out. This means that the washout applies to the Higgs asymmetry and the $N_{2}$-decay produced $\Delta_{\alpha}$ asymmetry, not to $\Delta_{\beta\neq\alpha}$. This is the primary source of flavour coupling. Secondarily, the redistribution of the asymmetries into quarks and RH charged particles also couples the $\Delta_{\alpha}$ asymmetries.
@@ 3.1 A more tangled expression for the final asymmetry
Solving the system of coupled differential equations at the production, for $T\sim M_{2}$, and at the $N_{1}$-washout, for $T\sim M_{1}$, leads to the following analytical expression for each final $\Delta_{\alpha}$ asymmetry fuller :
@@
Before explaining in detail the meaning and definition of all involved quantities in this general expression, notice that now the final BLB-L asymmetry, though still given by $N^{\rm f}_{B-L}=\sum_{\alpha=e,\mu,\tau}N^{\rm f}_{\Delta_{\alpha}}$, is the sum of 27 terms instead of just 3. The matrix
@@
is the matrix that diagonalizes $P^{0}_{1}$, i.e. $V\,P^{0}_{1}\,V^{-1}=P^{0}_{1^{\prime\prime}}\equiv{\rm diag}(P^{0}_{1e^{\prime\prime}},P^{0}_{1\mu^{\prime\prime}},P^{0}_{1\tau^{\prime\prime}})$, where we defined $P^{0}_{1\alpha}\equiv K_{1\alpha}/K_{1}$ and the rotated flavoured decay parameters as $K_{1\alpha^{\prime\prime}}\equiv P^{0}_{1\alpha^{\prime\prime}}\,K_{1}$. The matrix $P^{0}_{1}$ is given by
@@
and the three-flavour coupling matrix Davidson:2008bu
$C^{(3)}\equiv\left(\begin{array}[]{ccc}C_{ee}^{(3)}&C_{e\mu}^{(3)}&C_{e\tau}^{(3)}\\ C_{\mu e}^{(3)}&C_{\mu\mu}^{(3)}&C_{\mu\tau}^{(3)}\\ C_{\tau e}^{(3)}&C_{\tau\mu}^{(3)}&C_{\tau\tau}^{(3)}\end{array}\right)=\left(\begin{array}[]{ccc}188/179&32/179&32/179\\ 49/358&500/537&142/537\\ 49/358&142/537&500/537\end{array}\right)\,.$ (44)
@@
i.e. $U\,P^{0}_{2}\,U^{-1}={\rm diag}(P^{0}_{2\tau^{\bot^{\prime}}_{2}},P^{0}_{2\tau^{\prime}})$, where the two-flavour coupling matrix is given by Davidson:2008bu
@@
and we defined $P^{0}_{2\alpha}\equiv K_{2\alpha}/K_{2}$. We have now fully defined all quantities entering the expression for the flavoured asymmetries in Eq. (40). As we said, the final BLB-L asymmetry is still obtained as the sum of the three flavoured asymmetries and it can be checked that, for $U=V=I$, one recovers the expression Eq. (27) obtained neglecting flavour coupling. We can also unpack the first sum in Eq. (40) and write explicitly:
@@
This expression shows how now each flavoured asymmetry $N^{\rm f}_{\Delta_{\alpha}}$ is not simply given by one term containing the $N_{1}$ washout exponential suppression term described by $e^{-3\pi\,K_{1\alpha}/8}$ but it also contains terms that are washed out by exponentials $e^{-3\pi\,K_{1\delta^{\prime\prime}}/8}$ with $\delta\neq\alpha$. In this way, even though $K_{1\alpha}\gg 1$, there can still be unsuppressed contributions to $N^{\rm f}_{\Delta_{\alpha}}$ from terms with $K_{1\delta}\ll 1$. Even though these terms are weighted by factors $V^{-1}_{\alpha\delta}$, containing off-diagonal terms of the $C^{(3)}$ matrix that are ${\cal O}(0.1)$, they might be dominant in some cases and therefore, in general, they have to be accounted for.
@@ 3.2 Successful leptogenesis solutions
Despite the much higher intricacy of the expression Eq. (40), obtained accounting for flavour coupling, compared to Eq. (27), obtained neglecting flavour coupling, the discussion made in 2.3 on the parameter dependence and the consequent successful leptogenesis condition Eq. (39) still holds. Flavour coupling does not introduce any new parameter. It only introduces many new terms in the expression for the final asymmetry that are proportional to the off-diagonal terms in the flavour coupling matrix. These are usually sub-dominant and give a small correction but for particular values of the parameters they might yield new solutions in regions of parameter space that would otherwise be unaccessible when flavour coupling is neglected.
The search of these solutions is the main objective of including flavour coupling. The result is shown in the right panel of Fig. 1. Here we show again the three-dimensional projection of a scatter plot containing about $2\times 10^{6}$ solutions obtained for $\alpha_{2}=5$ and scanning over the 15 parameters in $m_{\nu}$ and $V_{L}$, respecting the experimental constraints in Eq. (6). The solutions have been again obtained imposing $\chi^{2}<\chi^{2}_{\rm max}=25$ and $\eta_{B}>6.01\times 10^{-10}$. The first thing one can notice is that the main structure of the allowed region is not changed by flavour coupling and, in particular, one still has a lower bound on $m_{1}$. On the other hand, it can be noticed how there are new muonic solution appearing in a region that was before excluded, at values of $m_{1}\sim 10\,{\rm meV}$. Moreover, electronic solutions, indicated by red points, now appear in the region of $\tau_{A}$ solutions.
@@ 3.2.1 Two new types of solutions
Let us have a closer look a these new solutions. As we did in the last section for the three types of old solutions, in Fig. 9 we plot different quantities, $M_{I},\varepsilon_{\alpha},K_{1\alpha},K_{2\tau}$ and $K_{2\tau^{\bot}_{2}}$, versus $m_{1}$ for two benchmark new solutions, one muonic and one electronic, with values of the parameters given in Table 2.
@@
As one can see, successful leptogenesis is satisfied for $m_{1}=42\,{\rm meV}$ and $m_{1}=2.1\,{\rm meV}$ for the muon-dominated and the electron-dominated solution, respectively. One can also see how the dominant CPC\!P asymmetry is the tauonic one, while both the muonic and the electronic CPC\!P asymmetries would be too small, in the absence of flavour coupling, to yield the observed baryon asymmetry. Therefore, the existence of these two new types of solutions crucially relies on the account of flavour coupling (without flavour coupling the muon and electron asymmetries, respectively, would be orders of magnitude below the observed value). We should add, however, that these new electronic solutions are very marginal ones. This can be somehow understood by the fact that the peak of the asymmetry in the shown benchmark case is just above the observed value. In the $2\times 10^{6}$ point scatter plot we generated, we obtained only less than $200$ electronic solution (i.e., less than $0.1\%$). The new muonic solutions are less special and we obtained about $2\times 10^{3}$ of them (i.e., $\sim 1\%$).
It is useful to give some analytical insight. Starting from Eq. (3.1), we can select in the second sum only terms with $\beta=\tau$ and focus only on the muonic and tauon asymmetry, if we consider the muonic benchmark solution in Fig. 9, obtaining:
$N^{\rm f}_{\Delta\mu}$ $=$ $N_{\Delta\tau}^{T\sim T_{L}}\left[V^{-1}_{\mu\mu^{\prime\prime}}V_{\mu^{\prime\prime}\tau}e^{-{3\pi\over 8}K_{1\mu^{\prime\prime}}}+V^{-1}_{\mu\tau^{\prime\prime}}V_{\tau^{\prime\prime}\tau}e^{-{3\pi\over 8}K_{1\tau^{\prime\prime}}}\right]$ (50)
$N^{\rm f}_{\Delta\tau}$ $=$ $N_{\Delta\tau}^{T\sim T_{L}}\left[V^{-1}_{\tau\mu^{\prime\prime}}V_{\mu^{\prime\prime}\tau}e^{-{3\pi\over 8}K_{1\mu^{\prime\prime}}}+V^{-1}_{\tau\tau^{\prime\prime}}V_{\tau^{\prime\prime}\tau}e^{-{3\pi\over 8}K_{1\tau^{\prime\prime}}}\right]\,.$ (51)
In the absence of flavour coupling, for $V=I$, one recovers $N^{\rm f}_{\Delta\mu}=0$, since we are neglecting terms $\propto N_{\Delta\mu}^{T\sim T_{L}}$, and $N^{\rm f}_{\Delta\tau}\simeq N_{\Delta\tau}^{T\sim T_{L}}e^{-{3\pi\over 8}K_{1\tau}}$. When flavour coupling is taken into account the muon asymmetry can get filled while the tauonic asymmetry can get depleted since, though the matrix V and its inverse contain negative terms. Notice that both in Eq. (50) and in Eq. (51) there is apparently no role played by the processes involving the electron flavour, the only one undergoing a strong wash-out ($K_{1e}^{\prime\prime}\gg 1)$ while the muon and tauon flavours seem to be decoupled ($K_{1\mu}^{\prime\prime}$, $K_{1\tau}^{\prime\prime}\lesssim 1$). One can wonder then why the $V$ and $V^{-1}$ entries coupling muon and tauon flavours should not be vanishing. The answer is that these entries are indeed sensitive to the large value of $K_{1e}$ resulting in large entries of the first row in the matrix $P^{0}_{1}$. Physically, this is because of the strong modification of the Higgs asymmetry due to electron washout that yields a strong feedback producing contribution to the the asymmetries in the muon and tauon flavour in a way that the muon asymmetry gets enhanced and the tauon asymmetry gets suppressed. [^15] As an example, we give here the matrices $P^{0}_{1}$, $V$ and $V^{-1}$ for the benchmark muon solution with values of the parameters in Table 2 well illustrating the role played by the large washout in the electron flavour: $P^{0}_{1}=\begin{pmatrix}1.039&0.177&0.177\\ 0.0007&0.0048&0.0014\\ 0.00076&0.0015&0.0052\end{pmatrix}\,,\;V=\begin{pmatrix}0.99999&0.00068&0.00073\\ -0.000054&0.77&-0.64\\ -0.001&0.64&0.77\end{pmatrix}\,,\;V^{-1}=\begin{pmatrix}0.99999&-0.00005&-0.000998\\ 0.00068&0.77&0.638\\ 0.00073&-0.64&0.77\end{pmatrix}\,$. (52)
@@ 3.2.2 Lower bound on $m_{1}$
In the right panel of Fig. 3 we again show the isocontour lines for the $m_{1}$ lower bound in the $\delta$ versus $\theta_{23}$ plane. One can see how they are just slightly modified by the inclusion of flavour coupling and, for example, the absolute minimum value gets slightly relaxed from $m_{1}\geq 0.89\,{\rm meV}$ to $m_{1}\geq 0.65\,{\rm meV}$. The most remarkable difference is that now the excluded region by the cosmological upper bound Eq. (2), in red in Fig. 3, gets very stronly reduced in a way that there are solutions compatible with the cosmological upper bound (blue region) basically for all allowed experimental values of $\delta$ and $\theta_{23}$, even at $3\sigma$.
@@ 3.2.3 Upper bound on $\theta_{23}$
In the right panel of Fig. 4, we show how flavour coupling affects the upper bound on $\theta_{23}$ versus $m_{1}$. One can see that the new muonic solutions tend to cover, very marginally, the region for $m_{1}\simeq 15\,{\rm meV}$ and $\theta_{23}\simeq 44^{\circ}$$45^{\circ}$ that would be excluded without an account of flavour coupling. However, all second octant values of $\theta_{23}$ still remain excluded for $m_{1}=(10$$30)\,{\rm meV}$.
@@ 3.2.4 Excluded regions for $\delta$
In the right panel of Fig. 5 we show the constraints in the plane $\delta$ versus $m_{1}$. One can see how the new muonic solutions tend to fill some existing gaps for $m_{1}\simeq 10\,{\rm meV}$ and $\delta$ in the range between $\pi/2$ and $3\pi/2$. However, the hole at $\delta\simeq\pi/2$, though shrinking, still survives. For $m_{1}$ much below 10 meV the constraints are practically unchanged.